The market for lemon has 10 potential consumers, each having an individual demand curve

P = 101 - 10Qi, where P is price in dollars per cup and Qi is the number of cups demanded

per week by the ith consumer. Find the market demand curve using algebra. Draw an

individual demand curve and the market demand curve. What is the quantity demanded by

each consumer and in the market as a whole when lemon is priced at P = $1/cup?


1
Expert's answer
2021-07-12T05:19:04-0400

If P = 101 - 10Qi, then Qi = 10.1 - 0.1P.

The market demand is:

Q = 10Qi = 101 - P.

If lemon is priced at P = $1/cup, then:

Q = 101 - 1 = 100 cups sold.